# Designing Speech Technologies for Australian Aboriginal English: Opportunities, Risks and Participation

### v1

@@ 2.1. Australian Aboriginal English
These negative attitudes are exacerbated because Aboriginal English is primarily oral, with no nationally adopted guidelines as to how it should be written. Additionally, the orthographic representation of Aboriginal English varies across locales, reflecting the local linguistic biographies and practices of distinct Indigenous groups. For example, the name of the Traditional Owners of the region comprising the southwest of Western Australia, as well as their language, is variably spelled ‘Nyungar’, ‘Noongar’, or ‘Nyoongar’ depending on community members’ preferences. [^1] Author 3 is Nyungar so we prefer this spelling. Rather than a weakness, the lack of standardising orthographies for Aboriginal English reflects Indigenous people’s prerogative to differentiate—rather than homogenise—how Aboriginal English is used across Indigenous Country.
Indigenous people tend to find the use of direct questions awkward or offensive (Rodríguez Louro and Collard, 2024), [^2] With some exceptions, including teaching contexts where questions instructors asking questions is heavily routinised (Moses et al., 2008). and these pragmatic considerations continue in Aboriginal English. Direct questions are also likely to elicit misleading information; Indigenous people have been known to answer ‘yes’ or agree with non-Indigenous people even if the question is not fully understood—a phenomenon known as ‘gratuitous concurrence’ (Liberman, 1985, p. 137). To avoid these limitations, researchers have used ‘yarning’, an Indigenous cultural form of storytelling and conversation, to lead unstructured data collection with Indigenous communities who speak Aboriginal English (Anonymous, YYYY).
@@ 2.4. Culture, and participatory methods, in language research
Within sociolinguistics, the ‘yarning’ method has demonstrated the effectiveness of culturally appropriate, participatory research methods when working with Aboriginal communities on speech data collection projects (Anonymous, YYYY). ‘Yarning’ favours group-based unstructured storytelling over individual interviews, which aligns with the Indigenous cultural protocol of centring the group over the individual (Malcolm, 2018, p. 17). Engaging in groups makes people feel less self-conscious in speech data collection activities than in one-on-one interactions with a researcher. Previous research led by Author 3, who is Indigenous, demonstrates how using such culturally appropriate, participatory methods throughout the research process—from choosing data collection protocols and fieldwork location to participant choice—enables recording of the authentic voices of people infrequently represented in academic research.
@@ 4. Designing ASR for Australian Aboriginal English
We now discuss a real-world partnership between company anonymized for peer review and university anonymized for peer review which aims to improve ASR for Australian Aboriginal English. We focus on the design considerations, and specifically on our efforts to relate the project goals and processes to both the sociolinguistic context (§2.1) and the technological opportunities (§3.1). We also discuss how we attempt to mitigate the system and project risks (§3.2, §3.4) while providing societal and economic opportunities (§3.3), e.g., by integrating participatory processes throughout the project lifecycle.
@@ 4.2. Elicitation of speech data
The collection protocols to secure the speech data needed for this project were designed by the authors with advice from the Indigenous Advisory Committee (§4.5). The project also applies participatory research methods discussed further in §7 and previously used by Author 2 and Author 3 (Anonymous, YYYY). However, the data to be collected for this project will be used by company anonymized for peer review and it is expected that the corpus will be open access, in accordance to FAIR and CARE principles (Wilkinson et al., 2016; Carroll et al., 2020), subject to further input from the Indigenous Advisory Committee. For this reason, the project explicitly excludes sharing of taboo cultural materials and sensitive information that is potentially harmful for Indigenous people and communities. To ensure cultural safety (cf. Risk 7), we designed a program of careful elicitation by using culturally relevant and sensitive prompts in consultation with the Indigenous Advisory Committee for this project, and ultimately designed by an Indigenous-owned company which specialises in producing educational resources using an anti-racism lens (Appendix A, cf. Risk 4).
@@ 4.3. Transcription of speech data
To co-develop transcription guidelines with community members, we required examples of public Aboriginal English to refer to. For this, we used a corpus of over 100 hours of YouTube videos that are primarily in Aboriginal English (see Appendix B). An Indigenous team, led by Author 2, developed the corpus and used their expertise with Aboriginal English to validate the variety spoken, ensuring that speakers were only coded as Aboriginal English if they used at least two linguistic features that distinguish this variety from Standardised Australian English. We developed an initial version of variety-specific transcription guidelines using examples from this corpus. Example guidelines include:
@@ 4.5. Reflections
We acknowledge the colonial context in which this project takes place. Aboriginal and Torres Strait Islander people, the traditional owners of Australia, represent 3.8% of the national population (Australian Government et al., 2020), and are minoritised by the non-Indigenous mainstream population, with significant gaps on economic, educational and health indicators, and life expectancy (Dockery and Milsom, 2007; Commonwealth of Australia, 2024; Australian Institute of Health and Welfare, 2023a, b). Most ancestral languages are extinct or endangered, and Aboriginal English is spoken by around 80% of Aboriginal and Torres Strait Islander people. This project focuses on both Aboriginal English varieties spoken in southwest Western Australia, where Author 2 and Author 3 live and work, and varieties of Aboriginal English spoken in the Pilbara (northern Western Australia), Queensland and Alice Springs (Northern Territory) (cf. Risk 2).
@@
The project is a partnership between company anonymized for peer review and university anonymized for peer review. company anonymized for peer review is a technology company, many of whose products incorporate ASR technologies. university anonymized for peer review is an Australian University whose Linguistics department has both expertise in Aboriginal English and existing relationships with Aboriginal communities, including the Nyungar community. Indigenous associates participate in the project, not just as mere providers of voice data, but as partners in the design of elicitation materials and transcription guidelines, management of the voice data collection, and validation and annotation of the collected materials (Opportunity 7). An Indigenous Advisory Committee was also established to provide guidance throughout the project. Investing in these relationships with partners who have earned community trust helps to mitigate risks to relationships with community (cf. Risk 5).
@@
The project is funded by company anonymized for peer review, who set the technological goal of the project: improving ASR for Aboriginal and Torres Strait people in Australia.
@@
Prior work by Author 1 and Author 4 has identified that improving ASR for Aboriginal English would bring practical benefits to Indigenous people in Australia (Anonymous, YYYY), by supporting spoken language in many applications (cf. Risk 1).
We recognise that the legacy of unethical and settler-colonial research has contributed to distrust of research in many Indigenous communities (Smith, 2021). This, along with a distrust of technology companies, can jeopardise even well intentioned projects. company anonymized for peer review found that seeking out partnership with (socio)linguists at university anonymized for peer review with existing relationships, cultural competencies, and sociolinguistic expertise, was critical, and preferable to attempting to build them from scratch (Doğruöz and Sitaram, 2022; Cooper et al., 2024).
@@ 4.7. Limitations
The technological goals of this project were set by company anonymized for peer review, and this precluded Indigenous people exercising full control over the project goals. We also acknowledge that Aboriginal and Torres Strait Islander people are less likely to be in full time research positions at companies or universities in Australia, limiting opportunities to be partner investigators on this research.
@@ 7. Ethical considerations
Before funding was secured for the project, Aboriginal scholars (with expertise in education, applied linguistics, and language revitalisation) were approached and asked to participate as members of a committee advising researchers on the project. Because the research involves Indigenous participants, letters of support from these advisors and from aiatsis (Australian Institute of Aboriginal and Torres Strait Islander Studies) were included in a submission to university anonymized for peer review’s IRB. Both university anonymized for peer review’s IRB and company anonymized for peer review’s privacy and legal teams reviewed the consent forms that would be used with Aboriginal people creating voice recordings.
@@
The collection of Aboriginal English voice recordings for the purposes of this project raises ICIP issues. As a team, the academic authors at university anonymized for peer review negotiated with company anonymized for peer review the inclusion of an ICIP clause in the research agreement signed in November 2023. This clause reads as follows:
@@
The collection of Aboriginal English for use by company anonymized for peer review must be approached with extreme care. From the start, Author 2 and Author 3 clarified with their partners that they would not use unstructured yarning sessions like they had done for Anonymous Grant Name (see, e.g., (Anonymous, YYYY)). With no pre-determined questions and clear Indigenous leadership through the research process, that project collected language as used in everyday settings, including sensitive stories about dispossession, trauma and death by suicide. These materials are firmly closed access with no risk of the audio files and associated transcriptions entering the public domain. However, unstructured yarning sessions tend to give rise to highly culturally sensitive material, which should be protected from use by third parties, especially large technology companies like company anonymized for peer review. For this reason, in consultation with Author 1 and the Advisory Committee, Author 2 and Author 3 recommended for this project the design of conversational prompts and story boards (see Appendix A) which seek to elicit interaction on a variety of general topics while steering clear of culturally sensitive material.
Once collected, all data will be screened by a team of Indigenous RAs, led by Author 3, and consisting of anonymised for peer review and anonymised for peer review, who will vet the audio files and transcriptions for cultural sensitivity. Any culturally unsafe material (e.g., stories of specific people, places, Indigenous knowledges, trauma, and loss) will be erased from the corpus before sharing the data with others.
@@
8. Not respecting Indigenous data sovereignty company anonymized for peer review not owning the collected speech data; Indigenous Advisory Committee guidance on appropriate Indigenous Data Governance.

### v2

@@ 2.1. Australian Aboriginal English
These negative attitudes are exacerbated because Aboriginal English is primarily oral, with no nationally adopted guidelines as to how it should be written. Additionally, the orthographic representation of Aboriginal English varies across locales, reflecting the local linguistic biographies and practices of distinct Indigenous groups. For example, the name of the Traditional Owners of the region comprising the southwest of Western Australia, as well as their language, is variably spelled ‘Nyungar’, ‘Noongar’, or ‘Nyoongar’ depending on community members’ preferences. [^1] Collard is Nyungar so we prefer this spelling. Rather than a weakness, the lack of standardising orthographies for Aboriginal English reflects Indigenous people’s prerogative to differentiate—rather than homogenise—how Aboriginal English is used across Indigenous Country.
Indigenous people tend to find the use of direct questions awkward or offensive (Rodríguez Louro and Collard, 2024), [^2] With some exceptions, including teaching contexts where questions instructors asking questions is heavily routinised (Moses et al., 2008). and these pragmatic considerations continue in Aboriginal English. Direct questions are also likely to elicit misleading information; Indigenous people have been known to answer ‘yes’ or agree with non-Indigenous people even if the question is not fully understood—a phenomenon known as ‘gratuitous concurrence’ (Liberman, 1985, p. 137). To avoid these limitations, researchers have used ‘yarning’, an Indigenous cultural form of storytelling and conversation, to lead unstructured data collection with Indigenous communities who speak Aboriginal English (rodriguez2021working).
@@ 2.4. Culture, and participatory methods, in language research
Within sociolinguistics, the ‘yarning’ method has demonstrated the effectiveness of culturally appropriate, participatory research methods when working with Aboriginal communities on speech data collection projects (Anonymous, YYYY). ‘Yarning’ favours group-based unstructured storytelling over individual interviews, which aligns with the Indigenous cultural protocol of centring the group over the individual (Malcolm, 2018, p. 17). Engaging in groups makes people feel less self-conscious in speech data collection activities than in one-on-one interactions with a researcher. Previous research led by Collard, who is Indigenous, demonstrates how using such culturally appropriate, participatory methods throughout the research process—from choosing data collection protocols and fieldwork location to participant choice—enables recording of the authentic voices of people infrequently represented in academic research.
@@ 4. Designing ASR for Australian Aboriginal English
We now discuss a real-world partnership between Google and University of Western Australia which aims to improve ASR for Australian Aboriginal English. We focus on the design considerations, and specifically on our efforts to relate the project goals and processes to both the sociolinguistic context (§2.1) and the technological opportunities (§3.1). We also discuss how we attempt to mitigate the system and project risks (§3.2, §3.4) while providing societal and economic opportunities (§3.3), e.g., by integrating participatory processes throughout the project lifecycle.
@@ 4.2. Elicitation of speech data
The collection protocols to secure the speech data needed for this project were designed by the authors with advice from the Indigenous Advisory Committee (§4.5). The project also applies participatory research methods discussed further in §7 and previously used by Rodríguez Louro and Collard (Rodríguez Louro and Collard, 2024). However, the data to be collected for this project will be used by Google and it is expected that the corpus will be open access, in accordance to FAIR and CARE principles (Wilkinson et al., 2016; Carroll et al., 2020), subject to further input from the Indigenous Advisory Committee. For this reason, the project explicitly excludes sharing of taboo cultural materials and sensitive information that is potentially harmful for Indigenous people and communities. To ensure cultural safety (cf. Risk 7), we designed a program of careful elicitation by using culturally relevant and sensitive prompts in consultation with the Indigenous Advisory Committee for this project, and ultimately designed by an Indigenous-owned company which specialises in producing educational resources using an anti-racism lens (Appendix A, cf. Risk 4).
@@ 4.3. Transcription of speech data
To co-develop transcription guidelines with community members, we required examples of public Aboriginal English to refer to. For this, we used a corpus of over 100 hours of YouTube videos that are primarily in Aboriginal English (see Appendix B). An Indigenous team, led by Rodríguez Louro, developed the corpus and used their expertise with Aboriginal English to validate the variety spoken, ensuring that speakers were only coded as Aboriginal English if they used at least two linguistic features that distinguish this variety from Standardised Australian English. We developed an initial version of variety-specific transcription guidelines using examples from this corpus. Example guidelines include:
@@ 4.5. Reflections
We acknowledge the colonial context in which this project takes place. Aboriginal and Torres Strait Islander people, the traditional owners of Australia, represent 3.8% of the national population (Australian Government et al., 2020), and are minoritised by the non-Indigenous mainstream population, with significant gaps on economic, educational and health indicators, and life expectancy (Dockery and Milsom, 2007; Commonwealth of Australia, 2024; Australian Institute of Health and Welfare, 2023a, b). Most ancestral languages are extinct or endangered, and Aboriginal English is spoken by around 80% of Aboriginal and Torres Strait Islander people. This project focuses on both Aboriginal English varieties spoken in southwest Western Australia, where Rodríguez Louro and Collard live and work, and varieties of Aboriginal English spoken in the Pilbara (northern Western Australia), Queensland and Alice Springs (Northern Territory) (cf. Risk 2).
@@
The project is a partnership between Google and University of Western Australia. Google is a technology company, many of whose products incorporate ASR technologies. University of Western Australia is an Australian University whose Linguistics department has both expertise in Aboriginal English and existing relationships with Aboriginal communities, including the Nyungar community. Indigenous associates participate in the project, not just as mere providers of voice data, but as partners in the design of elicitation materials and transcription guidelines, management of the voice data collection, and validation and annotation of the collected materials (Opportunity 7). An Indigenous Advisory Committee was also established to provide guidance throughout the project. Investing in these relationships with partners who have earned community trust helps to mitigate risks to relationships with community (cf. Risk 5).
@@
The project is funded by Google, who set the technological goal of the project: improving ASR for Aboriginal and Torres Strait people in Australia.
@@
Prior work by Hutchinson and Cooper has identified that improving ASR for Aboriginal English would bring practical benefits to Indigenous people in Australia (Cooper et al., 2024), by supporting spoken language in many applications (cf. Risk 1).
We recognise that the legacy of unethical and settler-colonial research has contributed to distrust of research in many Indigenous communities (Smith, 2021). This, along with a distrust of technology companies, can jeopardise even well intentioned projects. Google found that seeking out partnership with (socio)linguists at University of Western Australia with existing relationships, cultural competencies, and sociolinguistic expertise, was critical, and preferable to attempting to build them from scratch (Doğruöz and Sitaram, 2022; Cooper et al., 2024).
@@ 4.7. Limitations
The technological goals of this project were set by Google, and this precluded Indigenous people exercising full control over the project goals. We also acknowledge that Aboriginal and Torres Strait Islander people are less likely to be in full time research positions at companies or universities in Australia, limiting opportunities to be partner investigators on this research.
@@ 7. Ethical considerations
Before funding was secured for the project, Aboriginal scholars (with expertise in education, applied linguistics, and language revitalisation) were approached and asked to participate as members of a committee advising researchers on the project. Because the research involves Indigenous participants, letters of support from these advisors and from aiatsis (Australian Institute of Aboriginal and Torres Strait Islander Studies) were included in a submission to University of Western Australia’s IRB. Both University of Western Australia’s IRB and Google’s privacy and legal teams reviewed the consent forms that would be used with Aboriginal people creating voice recordings.
@@
The collection of Aboriginal English voice recordings for the purposes of this project raises ICIP issues. As a team, the academic authors at University of Western Australia negotiated with Google the inclusion of an ICIP clause in the research agreement signed in November 2023. This clause reads as follows:
@@
The collection of Aboriginal English for use by Google must be approached with extreme care. From the start, Rodríguez Louro and Collard clarified with their partners that they would not use unstructured yarning sessions like they had done for RA/4/20/4977 ‘Aboriginal English in the global city: Minorities and language change’ (see, e.g., (rodriguez2021working)). With no pre-determined questions and clear Indigenous leadership through the research process, that project collected language as used in everyday settings, including sensitive stories about dispossession, trauma and death by suicide. These materials are firmly closed access with no risk of the audio files and associated transcriptions entering the public domain. However, unstructured yarning sessions tend to give rise to highly culturally sensitive material, which should be protected from use by third parties, especially large technology companies like Google. For this reason, in consultation with Hutchinson and the Advisory Committee, Rodríguez Louro and Collard recommended for this project the design of conversational prompts and story boards (see Appendix A) which seek to elicit interaction on a variety of general topics while steering clear of culturally sensitive material.
Once collected, all data will be screened by a team of Indigenous RAs, led by Collard, and consisting of Narrier and Jones, who will vet the audio files and transcriptions for cultural sensitivity. Any culturally unsafe material (e.g., stories of specific people, places, Indigenous knowledges, trauma, and loss) will be erased from the corpus before sharing the data with others.
@@
8. Not respecting Indigenous data sovereignty Google not owning the collected speech data; Indigenous Advisory Committee guidance on appropriate Indigenous Data Governance.